use the substitution method to solve the following system of linear equations.
-3x - y = 3
-3x - 5y = -21
what is the solution?
(__,__)

Answer by checkley77(12569)

-(-3x - y = 3 ) multiply this equation by -1 & add.
-3x - 5y = -21
---------------------
-4y=-24
-4y=-24
y=-24/-4
y=6 ans.
-3x-6=3
-3x=3=6
-3x=9
x=9/-3
x=-3 ans.
Proof:
-3*-3-5*6=-21
9-30=-21
-21=-21
what is the solution?
(-3,6)